4,Toyota Corp.'s stock is $30 per share. Its expected return is 25% and variance is 14%. Honda Corp.'s stock is $18 per share. Its expected return is 20% and variance is 4%. Benz Corp.'s stock is $42 per share. Its expected return is 13% and variance 7%. What would be the expected return of a portfolio consisting of 50% Toyota and 50% Honda?

————%

* Place your answer in percentage form, say 5.99% and not .0599

5,toyota has an expected return of 23%, and a variance of 0.013. Honda has an expected return of 20%, and a variance of 0.006. The covariance between Toyota and Honda is 0.06. Using these data, calculate the variance of a portfolio consisting of 50% Toyota and 50% Honda.
Answer #1

4) Expected return = Wt*Rt + Wh*Rh

where, Wt and Wh = weights of Toyota and Honda

Rt and Rh = expected returns of Toyota and Honda

= 0.5*25 +0.5*20

= 12.5+10

= 22.5%

5)

variance = Wt2 * Vt +Wh2 * Vh + Wt * Wh * co-variance between Toyota and Honda

where, Vt and Vh = variance of Toyota and Honda

= 0.5^2 * 0.013 + 0.5^2 * 0.006 + 0.5 *0.5 * 0.06

variance of portfolio   = 0.01975